# D-69120 Heidelberg, Germany

### v1

@@ 6.2 Torque efficiency
Results of the efficiency $Q_{\Gamma}$ calculated with DDSCAT may possess positive (negative) values dependent on wavelength (see Fig. 4) corresponding to a direction of the angular velocity ${\omega}_{\mathrm{RAT}}$ being parallel (anti-parallel) to $\hat{e}_{1}$. Hence, different wavelength regimes create opposing torques on the grain aggregates. Subsequently, taking the sign change into account may considerably diminishing the net RAT for individual aggregates. In order to evaluate the average efficiency for different grain aggregates we both evaluate $\Gamma_{\mathrm{RAT}}$ over $\lambda$ (see Eq. 1) for each grain individually including cancelling effects of positive and negative spectral contribution of $Q_{\Gamma}(\lambda)$ as well as calculate $\Gamma_{\mathrm{RAT}}$ using the absolute value $|Q_{\Gamma}(\lambda)|$, which ignores the opposing spectral contributions to the net RAT. We find that averaging over the DDSCAT data with the absolute value of $|Q_{\Gamma}|$ results only in marginally higher RATs compared to taking the sign of $Q_{\Gamma}$ into account. We emphasize that the difference is insignificant compared to the variation between individual grain shapes. Hence, we utilize only the absolute value $|Q_{\Gamma}|$ in the following analysis, to make it possible to combine data of different grain shapes.
@@
6.3 Parameterization for the RAT efficiency
Since RATs depend on the wavelength $\lambda$ relative to the grain radius $a_{\mathrm{eff}}$ it is convenient to evaluate $|Q_{\Gamma}|$ as a function of $\lambda/a_{\mathrm{eff}}$ (see Fig. 6) instead of $\lambda$ alone. Here, the slopes of the different grain sizes overlap tightly, showing the self-similar nature that allowed the canonical parametrization (Eq. 2) for solid grains in the first place. Our results using grain aggregates shows a comparable constant part for $\lambda/a_{\mathrm{eff}}<1.8$, but do not follow the overall power-law parametrization for $\lambda/a_{\mathrm{eff}}>1.8$. Instead both materials SC and CA, reveal a steeper power-law that overlaps for all sizes from $\lambda/a_{\mathrm{eff}}\approx 3$ and then goes over into a shallower power-law for $\lambda/a_{\mathrm{eff}}\gtrsim 10$. A similar behavior of a broken power-law can be seen in previous studies e.g. in LH07 or for gaussian dust grains in Herranen et al. (2019), but is much more apparent here. Consequently, we suggest a modified parameterization for ballistic grain aggregates with two connected consecutive power-laws. Least square fits for the constant part (called $c$ in this study) and the two power-laws are performed for all data points of the grains sharing the same material and monomer-density. The fits of $c$ are performed for individual grain sizes in the interval $\lambda/a_{\mathrm{eff}}<3$. All sizes are fitted jointly for the first and second power-law with exponents $q_{1}$ and $q_{2}$ in the intervals $123<(\lambda/a_{\mathrm{eff}})<$ 12 and $(\lambda/a_{\mathrm{eff}})>20$. In practice the power-laws are fitted as linear functions to the logarithm of the data.
@@ 6.4 Net torques and angular velocities
The exact difference in RAT using the canonical or the modified parameterization and, subsequently, the magnitude of the angular velocity of the grains cannot a priori be quantified due to the dependency of RATs on the spectral characteristics of the radiation field. Therefore we calculate the maximal angular velocity $\omega_{\mathrm{RAT}}$ induced by RATs using both parameterizations and compare the results to those we get using the DDSCAT data of individual grains.
If the torque efficiency $Q_{\Gamma}(\lambda)$ for a single grain exhibits some sign changes over wavelength (see Fig. 4), this leads to a lower torque overall as different wavelengths exert RAT components in opposite directions. As mentioned above, the effect of positive and negative components of $Q_{\Gamma}$ cancelling out appears to be small compared to the overall variation in the strength of $Q_{\Gamma}$ between different grains. The torque along $\hat{a}_{1}$ is calculated for each grain following Eq. 1, where $Q_{\Gamma}$ is used without taking the absolute value. Then, the maximal angular velocity is calculated with Eq. 3, where $\tau_{\mathrm{drag}}$ and $I_{\mathrm{max}}$ are also evaluated individually for each grain.
Figure 8 shows the distribution of angular velocities and torques for individual CA/BA grains and the most luminous source i.e. the AGN. The distributions are broken down by grain size. We report that the resulting values of the RATs may span a range up to two orders of magnitude, with a larger spread below the mean. The mean of the distribution increases with grains size, as larger grains have higher $Q_{\Gamma}$ values in the constant part. The distributions of the different grain sizes are more distinct in $\Gamma_{\mathrm{RAT}}$ and overlap significantly for $\omega_{\mathrm{RAT}}$, due to the influence of the drag timescale (see Eq. 3).
@@
Figure 8: Histograms representing the distribution of the RAT $\Gamma_{\mathrm{RAT}}$ (left) and angular velocity $\omega_{\mathrm{RAT}}$ (right) for CA/BA grains of different grain sizes showing and the AGN as radiation source. Continuous lines show the mean for each distribution, dotted lines show the value calculated using the canonical parametrization (LH07). Mean values for $\omega_{\mathrm{RAT}}$ are calculated from the mean values of $\Gamma_{\mathrm{RAT}}$ shown in the lower panel as well as averages for $\tau_{\mathrm{drag}}$ and $I_{\mathrm{max}}$. We note that RATs by the canonical parametrization are consistently higher than the mean values of the dust aggregates.
@@
To quantify how the canonical parametrization of LH07 and the modified parametrization of our study compare with the distribution means for all combinations of radiation sources, materials, and monomer-densities, we show the distribution means of the RATs in Fig. 9 as a function the grain size $a_{\mathrm{eff}}$. The CA material in the left column, the SC material in the right column, the monomer-densities in the rows and the four radiation sources are shown with different colours in each panel. The difference between the sources is due to the linear dependence of $\Gamma_{\mathrm{RAT}}$ with energy density $u$ and anisotropy factor $\gamma$ where more efficient RATs are reached for stronger radiation sources (see Eq. 1 and Tab. 2).
@@
In Fig. 10 we present the mean angular velocities $\omega_{\mathrm{RAT}}$ corresponding to the RATs shown in Fig. 9. Here, the variations of $\omega_{\mathrm{RAT}}$ between individual sources is different to the one of $\Gamma_{\mathrm{RAT}}$, because $\omega_{\mathrm{RAT}}$ also depends on $\tau_{\mathrm{drag}}$, which itself depends indirectly on $u$ via the temperature dependence of $\tau_{\mathrm{IR}}$. Using the dust temperatures in Tab. 3 one can estimate $\tau_{\mathrm{IR}}$ for the given temperature from Fig. 3. The gas drag timescale is independent of $u$ and about $\tau_{\mathrm{gas}}\gtrsim 10^{12}\,\mathrm{s}$ for the chosen gas density of $n_{\mathrm{gas}}=5\cdot 10^{7}\,\mathrm{m^{-3}}$. We can see, that the drag timescale is dominated by $\tau_{\mathrm{gas}}$ for the low dust temperatures of the ISRF source, both timescales are about equal for S1 and $\tau_{\mathrm{IR}}$ dominates for S2 and AGN, where the dust is the hottest.
@@
Therefore, the difference in $\omega_{\mathrm{RAT}}$ depends only on $u$ and $\gamma$ for the ISRF, but is dominated by $\tau_{\mathrm{IR}}$ for the other radiation sources. This is why the three stronger sources are much closer in $\omega_{\mathrm{RAT}}$ than in $\Gamma_{\mathrm{RAT}}$. The difference in $\tau_{\mathrm{IR}}$ between the two materials thereby also leads to smaller angular velocity for S1, S2 and AGN for SC compared to CA, but no difference between the two materials for ISRF, as $\tau_{\mathrm{gas}}$ dominates there. The difference between monomer-densities is a decrease in $\Gamma_{\mathrm{RAT}}$ from BA to BAM2 smaller than one order of magnitude, due to $Q_{\Gamma}$ decreasing with monomer-density. The resulting decrease with monomer-density in $\omega_{\mathrm{RAT}}$ is smaller, due to $I_{\mathrm{max}}$ also decreasing with monomer-density.
As expected RATs have a strong scaling with $a_{\mathrm{eff}}$ and it does not change much between sources, materials and monomer-densities. However, for the angular velocity $\omega_{\mathrm{RAT}}$, the results are less uniform as shown in Fig. 10. Here, there is no scaling with grain size for the ISRF source, and a weak one for the other sources, which flattens towards the larger grains for SC and even declines for CA with the S1 source. We remark that the scaling of $\omega_{\mathrm{RAT}}$ with grain size depends on the interplay between multiple factors: spectrum of the source, magnitude of the drag timescale, which depends on the material and size of the grain, and whether $\tau_{\mathrm{IR}}$ or $\tau_{\mathrm{gas}}$ dominates for the given environment. It thereby differs from the scaling in other publications e.g. Hoang et al. (2019).
@@
The canonical parametrization gives more accurate results for the larger grains, as they experience the strongest torques. We point out that overall the canonical parametrization overestimates the mean $\omega_{\mathrm{RAT}}$, with differences up to one order of magnitude. In comparison, the two-power-law parametrization resembles the mean of the DDSCAT data more precisely and performs equally well for all sources, materials, monomer-densities, and sizes.
@@
|${Q}_{\Gamma}|=\begin{cases}0.4&\mbox{if }\lambda\leq 1.8a_{\mathrm{eff}}\times\phi\,,\\ 0.4\left(\frac{\lambda}{1.8a_{\mathrm{eff}}\ \times\ \phi}\right)^{\alpha}&\mbox{otherwise}\,.\end{cases}$ (14)
Fig. 11 shows the average angular velocity $\omega_{\mathrm{RAT}}$ in comparison with the TK21 parametrization and canonical parametrization of LH07. The proposed parametrization of TK21 tends to overestimate the resulting $\omega_{\mathrm{RAT}}$ as well, especially for the stronger radiation sources, but clearly improves upon the canonical parametrization of LH07. As shown in Fig. 11, merely introducing $\phi$ to take the grain porosity into account cannot improve the agreement between our results and the canonical parametrization considerably indicating that complex grain shapes differ in their interaction with radiation compared to compact grains in more fundamental ways than just by making the grains porous.
Consequently, using the two-power-law parametrization of $Q_{\Gamma}$ to calculate the average $\omega_{\mathrm{RAT}}$ for an ensemble of ballistic grain aggregates is the most accurate choice even though individual grains may have larger or much smaller $\omega_{\mathrm{RAT}}$ by several orders of magnitude (see Fig. 8), due to random variation in grain shape.
Figure 9: The RATs $\Gamma_{\mathrm{RAT}}$ (solid line) per grain size $a_{\mathrm{eff}}$ calculated with the two-power-law parametrization is shown in comparison to the RATs predicted by the canonical parametrization suggested in LH07 (dotted line) as well as the average of the DDSCAT simulation data (diamonds). Each color represents one of the different sources i.e. ISRF (grey), S1 (dark blue), S2 (light blue), or AGN (red), respectively, utilizing the CA grain material (left column) and the SC material (right column) with monomer-densities BA (top row), BAM (middle row), or BAM2 (bottom row). This figure is to be compared with Fig. 8.
Figure 10: Same as Fig. 9 but for angular velocity $\omega_{\mathrm{RAT}}$. We emphasize that the canonical parameterization may overpredict the magnitude of $\omega_{\mathrm{RAT}}$ up to one order of magnitude higher for BAM grain aggregates as the two-power-law parameterization.
Figure 11: The same as Fig. 10 but only for CA/BAM2 (left panel) and SC/BAM2 (right panel) grain aggregates showing the canonical paramterization of LH07 (dotted lines) in comparison with the paramterization for porous grains suggested in TK21 (dashed lines).

### v2

@@ 6.2 Torque efficiency
Results of the efficiency $Q_{\Gamma}$ calculated with DDSCAT may possess positive (negative) values dependent on wavelength (see Fig. 4) corresponding to a direction of the angular velocity ${\omega}_{\mathrm{RAT}}$ being parallel (anti-parallel) to $\hat{e}_{1}$. Hence, different wavelength regimes create opposing torques on the grain aggregates. Subsequently, taking the sign change into account may considerably diminishing the net RAT for individual aggregates. In order to evaluate the average efficiency for different grain aggregates we both evaluate $\Gamma_{\mathrm{RAT}}$ over $\lambda$ (see Eq. 1) for each grain individually including canceling effects of positive and negative spectral contribution of $Q_{\Gamma}(\lambda)$ as well as calculate $\Gamma_{\mathrm{RAT}}$ using the absolute value $|Q_{\Gamma}(\lambda)|$, which ignores the opposing spectral contributions to the net RAT. We find that averaging over the DDSCAT data with the absolute value of $|Q_{\Gamma}|$ results only in marginally higher RATs compared to taking the sign of $Q_{\Gamma}$ into account. We emphasize that the difference is insignificant compared to the variation between individual grain shapes. Hence, we utilize only the absolute value $|Q_{\Gamma}|$ in the following analysis, to make it possible to combine data of different grain shapes.
@@
6.3 Parametrization for the RAT efficiency
Since RATs depend on the wavelength $\lambda$ relative to the grain radius $a_{\mathrm{eff}}$ it is convenient to evaluate $|Q_{\Gamma}|$ as a function of $\lambda/a_{\mathrm{eff}}$ (see Fig. 6) instead of $\lambda$ alone. Here, the slopes of the different grain sizes overlap tightly, showing the self-similar nature that allowed the canonical parametrization (Eq. 2) for solid grains in the first place. Our results using grain aggregates shows a comparable constant part for $\lambda/a_{\mathrm{eff}}<1.8$, but do not follow the overall power-law parametrization for $\lambda/a_{\mathrm{eff}}>1.8$. Instead both materials SC and CA, reveal a steeper power-law that overlaps for all sizes from $\lambda/a_{\mathrm{eff}}\approx 3$ and then goes over into a shallower power-law for $\lambda/a_{\mathrm{eff}}\gtrsim 10$. A similar behavior of a broken power-law can be seen in previous studies e.g. in LH07 or for gaussian dust grains in Herranen et al. (2019), but is much more apparent here. Consequently, we suggest a modified parametrization for ballistic grain aggregates with two connected consecutive power-laws. Least square fits for the constant part (called $c$ in this study) and the two power-laws are performed for all data points of the grains sharing the same material and monomer-density. The fits of $c$ are performed for individual grain sizes in the interval $\lambda/a_{\mathrm{eff}}<3$. All sizes are fitted jointly for the first and second power-law with exponents $q_{1}$ and $q_{2}$ in the intervals $123<(\lambda/a_{\mathrm{eff}})<$ 12 and $(\lambda/a_{\mathrm{eff}})>20$. In practice the power-laws are fitted as linear functions to the logarithm of the data.
@@ 6.4 Net torques and angular velocities
The exact difference in RAT using the canonical or the modified parametrization and, subsequently, the magnitude of the angular velocity of the grains cannot a priori be quantified due to the dependency of RATs on the spectral characteristics of the radiation field. Therefore we calculate the maximal angular velocity $\omega_{\mathrm{RAT}}$ induced by RATs using both parametrizations and compare the results to those we get using the DDSCAT data of individual grains.
If the torque efficiency $Q_{\Gamma}(\lambda)$ for a single grain exhibits some sign changes over wavelength (see Fig. 4), this leads to a lower torque overall as different wavelengths exert RAT components in opposite directions. As mentioned above, the effect of positive and negative components of $Q_{\Gamma}$ canceling out appears to be small compared to the overall variation in the strength of $Q_{\Gamma}$ between different grains. The torque along $\hat{a}_{1}$ is calculated for each grain following Eq. 1, where $Q_{\Gamma}$ is used without taking the absolute value. Then, the maximal angular velocity is calculated with Eq. 3, where $\tau_{\mathrm{drag}}$ and $I_{\mathrm{max}}$ are also evaluated individually for each grain.
Figure 8 shows the distribution of angular velocities and torques for individual SC/BA grains and the most luminous source i.e. the AGN. The distributions are broken down by grain size. We report that the resulting values of the RATs may span a range up to two orders of magnitude, with a larger spread below the mean. The mean of the distribution increases with grain size, as larger grains have higher $Q_{\Gamma}$ values in the constant part. The distributions of the different grain sizes are more distinct in $\Gamma_{\mathrm{RAT}}$ and overlap significantly for $\omega_{\mathrm{RAT}}$, due to the influence of the drag timescale (see Eq. 3).
@@
Figure 8: Histograms representing the distribution of the RAT $\Gamma_{\mathrm{RAT}}$ (left) and angular velocity $\omega_{\mathrm{RAT}}$ (right) for SC/BA grains of different grain sizes and the AGN as radiation source. Continuous lines show the mean for each distribution, dotted lines show the value calculated using the canonical parametrization (LH07). Mean values for $\omega_{\mathrm{RAT}}$ are calculated from the mean values of $\Gamma_{\mathrm{RAT}}$ shown in the lower panel as well as averages for $\tau_{\mathrm{drag}}$ and $I_{\mathrm{max}}$. We note that RATs by the canonical parametrization are consistently higher than the mean values of the dust aggregates.
@@
To quantify how the canonical parametrization of LH07 and the modified parametrization of our study compare with the distribution means for all combinations of radiation sources, materials, and monomer-densities, we show the distribution means of the RATs in Fig. 9 as a function the grain size $a_{\mathrm{eff}}$. The CA material in the left column, the SC material in the right column, the monomer-densities in the rows and the four radiation sources are shown with different colors in each panel. The difference between the sources is due to the linear dependence of $\Gamma_{\mathrm{RAT}}$ with energy density $u$ and anisotropy factor $\gamma$ where larger RATs are reached for stronger radiation sources (see Eq. 1 and Tab. 2). The RAT of the S1 source should be larger by a factor of 7 due to $\gamma$ and by a factor of $5.6$ due to $u$. The difference seen in Fig. 9 is larger than that with approximately 2 orders of magnitude. This is a result of the spectral energy distribution of this source. The integral in Eq. 1 puts a much larger weight on $u$ at wavelengths smaller than $\approx 1\,\mu\mathrm{m}$, as $Q_{\lambda}$ drops by about 4 orders of magnitude for larger wavelengths. This is the range where the stellar sources outshine the ISRF.
@@
In Fig. 10 we present the mean angular velocities $\omega_{\mathrm{RAT}}$ corresponding to the RATs shown in Fig. 9. Here, the variations of $\omega_{\mathrm{RAT}}$ between individual sources is different to the one of $\Gamma_{\mathrm{RAT}}$, because $\omega_{\mathrm{RAT}}$ also depends on $\tau_{\mathrm{drag}}$, which itself depends indirectly on $u$ via the temperature dependence of $\tau_{\mathrm{IR}}$. Using the dust temperatures in Tab. 3 one can estimate $\tau_{\mathrm{IR}}$ for the given temperature from Fig. 3. The gas drag timescale is independent of $u$ and about $\tau_{\mathrm{gas}}\gtrsim 10^{12}\,\mathrm{s}$ for the chosen gas density of $n_{\mathrm{gas}}=5\cdot 10^{7}\,\mathrm{m^{-3}}$. We can see, that the drag timescale is dominated by $\tau_{\mathrm{gas}}$ for the lower dust temperatures of the ISRF and S1 sources, both timescales are about equal for S2 and $\tau_{\mathrm{IR}}$ dominates for AGN, where the dust is the hottest.
Figure 9: The RATs $\Gamma_{\mathrm{RAT}}$ (solid line) per grain size $a_{\mathrm{eff}}$ calculated with the two-power-law parametrization is shown in comparison to the RATs predicted by the canonical parametrization suggested in LH07 (dotted line) as well as the average of the DDSCAT simulation data (diamonds). Each color represents one of the different sources i.e. ISRF (gray), S1 (dark blue), S2 (light blue), or AGN (red), respectively, utilizing the CA grain material (left column) and the SC material (right column) with monomer-densities BA (top row), BAM (middle row), or BAM2 (bottom row). This figure is to be compared with Fig. 8.
Figure 10: Same as Fig. 9 but for angular velocity $\omega_{\mathrm{RAT}}$. We emphasize that the canonical parametrization may over-predict the magnitude of $\omega_{\mathrm{RAT}}$ up to one order of magnitude higher for BAM grain aggregates as the two-power-law parametrization.
@@
Therefore, the difference in $\omega_{\mathrm{RAT}}$ depends only on $\Gamma_{\mathrm{RAT}}$ for the ISRF and S1, but includes $\tau_{\mathrm{IR}}$ for S2 and AGN. This is why AGN is much closer to S2 in $\omega_{\mathrm{RAT}}$ ($1\,\mathrm{dex}$) than in $\Gamma_{\mathrm{RAT}}$ ($3\,\mathrm{dex}$). Even though the torque is much larger for the AGN source, the grains do not rotate that much faster, because they also get much hotter, increasing the IR-drag. The difference in $\tau_{\mathrm{IR}}$ between the two materials also leads to smaller angular velocity for S2 and AGN for SC compared to CA, but no difference between the two materials for S1 and ISRF, as $\tau_{\mathrm{gas}}$ dominates there. The difference between monomer-densities is a decrease in $\Gamma_{\mathrm{RAT}}$ from BA to BAM2 smaller than one order of magnitude, due to $Q_{\Gamma}$ decreasing with monomer-density. The resulting decrease with monomer-density in $\omega_{\mathrm{RAT}}$ is smaller, due to $I_{\mathrm{max}}$ also decreasing with monomer-density.
As expected RATs have a strong scaling with $a_{\mathrm{eff}}$ and it does not change much between sources, materials and monomer-densities. However, for the angular velocity $\omega_{\mathrm{RAT}}$, the results are less uniform as shown in Fig. 10. Here, there is no scaling with grain size for the ISRF source, and a weak one for the AGN, which flattens towards the larger grains. For the stellar sources $\omega_{\mathrm{RAT}}$ first increases and then flattens for SC, even decreasing towards larger grains for S1. In the case of CA $\omega_{\mathrm{RAT}}$ even scales inversely with grain size for the stellar sources. We remark that the scaling of $\omega_{\mathrm{RAT}}$ with grain size depends on the interplay between multiple factors: spectral shape of the source, scaling of the drag timescale, which depends on the material and structure of the grain, and whether $\tau_{\mathrm{IR}}$ or $\tau_{\mathrm{gas}}$ dominates for the given parameters. As outlined in the pioneering work of Hoang et al. (2019) and Hoang (2019) RATs may spin-up dust grains up to a characteristic angular velocity $\omega_{\mathrm{disr}}$ where the internal grain structure can no longer resist the disruption because of centrifugal forces (RATD). The predicted order of magnitude of for RATD to occur is $\omega_{\mathrm{disr}}\approx 10^{9}\mathrm{s}^{-1}$. Later, ab-initio numerical N-body simulations of the disruption of rotating BA aggregates revealed that aggregates larger than $a_{\mathrm{eff}}=300\ \mathrm{nm}$ are more resistant against RATD because of deformation. We emphasize that the most luminous sources, AGN and S2, respectively, may spin up the dust aggregates beyond the limit of $\omega_{\mathrm{RAT}}>\omega_{\mathrm{disr}}$ (see Fig. 10) even though RATs are less efficient for ballistic aggregates. Hence, most of the grains may become disrupted by RATD in proximity of an AGN or S2-like star. So ballistic aggregates are more likely to grow deep inside of dense molecular clouds where they are shielded from a strong radiation field (Bethell et al. 2007).
@@
As presented in Fig. 10, the canonical parametrization gives more accurate results for the larger grain aggregates, as they experience the strongest torques. We point out that overall the canonical parametrization overestimates the mean angular velocity $\omega_{\mathrm{RAT}}$, with differences up to one order of magnitude. In comparison, the two-power-law parametrization resembles the mean of the DDSCAT data more precisely and performs equally well for all sources, materials, monomer-densities, and sizes.
@@
|${Q}_{\Gamma}|=\begin{cases}0.4&\mbox{if }\lambda\leq 1.8\,a_{\mathrm{eff}}\cdot\phi\,,\\ 0.4\left(\frac{\lambda}{1.8\,a_{\mathrm{eff}}\,\cdot\,\phi}\right)^{\alpha}&\mbox{otherwise}\,.\end{cases}$ (14)
Fig. 11 shows the average angular velocity $\omega_{\mathrm{RAT}}$ in comparison with the TK21 parametrization and canonical parametrization of LH07. The model proposed by TK21 clearly improves upon the canonical parametrization of LH07 for the less porous BAM1 and BAM2 grains, but now slightly under-predicts $\omega_{\mathrm{RAT}}$ for the weaker sources in BAM1. For the more porous BA grains however TK21 clearly under-predicts $\omega_{\mathrm{RAT}}$, as this is the regime, where the canonical parametrization performed best. As shown in Fig. 11, merely introducing $\phi$ to take the grain porosity into account does not bring the canonical parametrization into agreement with our results, indicating that complex grain shapes differ in their interaction with radiation compared to compact grains in more fundamental ways than just by making the grains porous.
Consequently we may summarize, that even though there is a large spread in $\omega_{\mathrm{RAT}}$ between individual grains (see Fig. 8) due to random variation in grain shape, using a parametrization of $Q_{\Gamma}$ is still valid to model the average $\omega_{\mathrm{RAT}}$ for an ensemble of ballistic grain aggregates.
For the three parametrizations shown in this work the two-power-law parametrization is the most accurate choice for this task.
Figure 11: The same as Fig. 10 but showing the canonical parametrization of LH07 (dotted lines) in comparison with the parametrization for porous grains suggested in TK21 (dashed lines).